OpenStudy (anonymous):

integrands of pi/3 and 0 7 sin ^2 x/[sqrt. 1-cosx] dx

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/3} 7\sin^2x /[ \sqrt{1-cosx}] dx\]